40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14524     Entry time: Mon Apr 8 23:52:09 2019     In reply to: 14523     Reply to this: 14525   14527
Author: gautam 
Type: Update 
Category: ALS 
Subject: EX Green PDH error monitor calibrated 

Some time ago, I had done an actuator calibration of ITMX. This suspension hasn't been victim to the recent spate of suspension problems, so I can believe that the results of those measurement are still valid. So I decided to calibrate the in-loop error signal of the EX green PDH loop, which is recorded via an SR560, G=10, by driving a line in ITMY position (thereby modulating the X arm cavity length) while the EX green frequency was locked to the arm cavity length. Knowing the amount I'm modulating the cavity length by (500 cts amplitude sine wave at 33.14159 Hz using awggui, translating to ~17.2 kHz amplitude in green frequency), I demodulated the response in C1:ALS-X_ERR_MON_OUT_DQ channel. At this frequency of ~33 Hz, the servo gain should be large, and so the green laser frequency should track the cavity length nearly perfectly (with transfer function 1/(1+L), where L is the OLG).

The response had amplitude 5.68 +/- 0.10 cts, see Attachment #1. There was a sneaky gain of 0.86 in the filter module, which I saw no reason to keep at this strange value, and so updated to 1, correcting the demodulated response to 6.6 cts. After accounting for this adjustment, the x10 gain of the SR560, and the loop suppression, I put a "cts2Hz" filter in (Attachment #2). I had to guess a value for the OLG at 33 Hz in order to account for the in-loop suppression. So I measured the OLTF using the usual IN1/IN2 method (Attachment #3), and then used a LISO model of the electronics, along with guesses of the cavity pole (18.5 kHz), low-pass filter poles (4x real poles at 70 kHz), PZT actuator gain (1.7 MHz/V) and PDH discriminant (40 uV/Hz, see this elog) to construct a model OLTF. Then I fudged the overall gain to get the model to line up with the measurement between 1-10kHz. Per this model, I should have ~75dB of gain at ~33Hz, so the tracking error to my cavity length modulation should be ~3.05 Hz. Lines up pretty well with the measured value of 4.7 Hz considering the number of guessed parameters. The measured OLG tapers off towards low frequency probably because the increased loop suppression drives one of the measured inputs on the SR785 into the instrument noise floor.

The final calibration number is 7.1 Hz/ct, though the error on this number is large ~30%. Note that these "Hertz" are green frequency changes - so the change to the IR frequency will be half.

Attachment #4 shows the error signal in various conditions, labelled in the legend. Interpretations to follow.

Attachment 1: errMonCalib.pdf  19 kB  Uploaded Tue Apr 9 19:21:27 2019  | Hide | Hide all | Show all
errMonCalib.pdf
Attachment 2: errMon.png  22 kB  Uploaded Tue Apr 9 19:23:20 2019  | Hide | Hide all | Show all
errMon.png
Attachment 3: OLTF.pdf  215 kB  Uploaded Tue Apr 9 19:40:41 2019  | Show | Hide all | Show all
Attachment 4: EX_frequencyNoises.pdf  45 kB  Uploaded Tue Apr 9 19:42:15 2019  | Hide | Hide all | Show all
EX_frequencyNoises.pdf
ELOG V3.1.3-